OpenStudy (abbycross167):

I WILL FAN AND GIVE A MEDAL!! Can someone please help me? My teacher said that need to redo this problem but I don’t know how… I’ll add the Question then My Answer then My Teachers response.

OpenStudy (abbycross167):

Find the point-slope form of the equation that goes through these two points. (1, 20) and (8, 4.5)

OpenStudy (abbycross167):

My Answer~ Y2-y1/x2-x1 4.5-20/8-1 -15.5/7 =2.21428571429 =2.21

OpenStudy (mathmale):

Let's focus on that first. You're given 2 points, and are to find the SLOPE of the line that goes thru them.

OpenStudy (baru):

u seem to have dropped the negative sign...

OpenStudy (abbycross167):

What my teacher said~ "not finished"

OpenStudy (mathmale):

Here's some feedback on your work: That slope formula should be labelled: \[y=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (mathmale):

... and "m" would be a better label than "y" because "m" represents slope.

OpenStudy (mathmale):

Next, your 4.5-20 / 8-1 is subject to misinterpretation unless you use parentheses around both numerator and denominator: correct: m = (4.5-20) / (8-1) not correct: 4.5-20 / 8-1

OpenStudy (mathmale):

Your -15.5 / 7 is ok as typewritten by you. I'd suggest NOT finding the decimal equivalent. Instead, look for a way to reduce -15.5/7 to simpler form. Could you do that now, please?

OpenStudy (mathmale):

Hint: 15.5=31/2

OpenStudy (abbycross167):

I'm not sure how to, I'm sorry i'm really not good with math

OpenStudy (mathmale):

For your own self-image, please do not tell others (or yourself) that you're "really not good in math." Instead, you might say, "math has been a bit of a challenge for me, but with effort and practice, I can learn what I need to know about it." -15.5/7 could be re-written as (-15.5) / 7 and the -15.5 could be re-written as -31/2 So, inside the parentheses you have -31/2: (-31/2) / 7, or (easier to read):\[\frac{ \frac{ -31 }{ 2 } }{ 7 }\]

OpenStudy (mathmale):

can you simplify that? Hint: re-write the 7 as 7/1. Then divide:\[\frac{ \frac{ -31 }{ 2 } }{ \frac{ 7 }{ 1 } }\]

OpenStudy (abbycross167):

-2 3/14?

OpenStudy (mathmale):

Hints: Leave the -31/2 as is. Invert the 7/1 in the denominator and multiply the result by -31/2.

OpenStudy (mathmale):

Your result is correct. It'd be a bit easier to leave it as -31/14.

OpenStudy (mathmale):

Personally, I'd leave your answer in that latest form, and not convert that to a decimal fraction.

OpenStudy (mathmale):

OK with this? Next problem?

OpenStudy (abbycross167):

could you give me a second so I can look through the steps we just did?

OpenStudy (mathmale):

sure. I need a bit of time to make a phone call. return when you're ready.

OpenStudy (abbycross167):

So I re-wrote the steps that we just did, would you mind looking at them to see if I wrote them right?

OpenStudy (mathmale):

OK. type them in as quickly as you can. Thx.

OpenStudy (abbycross167):

Find the point-slope form of the equation that goes through these two points. (1, 20) and (8, 4.5) My Answer~ Y= y2-y1/x2-x1 (4.5-20)/(8-1) = -15.5/7 Rewrite -15.5/7 as (-15.5)/7 Rewrite -15.5 as -31/2 = -31/2 7/1 Divide: -31/2 7/1 = -31/14

OpenStudy (mathmale):

Please enclose your " -31/2 " and (separately ) your "7/1" in parentheses and show that you are dividing (-31/2) by (7/1). Currently your work doesn't show the actual division symbol, even though you do use the word "divide."

OpenStudy (abbycross167):

so like this? (-31/2) (7/1)

OpenStudy (mathmale):

Yes, except put the division symbol right after (-31/2).

OpenStudy (abbycross167):

ok

OpenStudy (mathmale):

Gotcha. I need to get off the 'Net, unfortunately, but just might be on O. S. again later today. Glad I got to work with you, and I appreciate your interest and perseverance.

OpenStudy (abbycross167):

umm ok

OpenStudy (abbycross167):

thank you for your help